Leapfrog Mode Testing and Observing Progress Log
Bold entries highlights filenames.
Purple text highlights data which is useable,
blue text highlights data containing measurements of the cals.
Red text highlights adjustments to integration times.


DSS image of Galaxy cluster group, partially observed on 18-19th nov with leapfrog
colours indicate a 'set' of leaps.



Date
Data Files written, Comments
General Comments
16th Nov
wapp.20041116.a1961.0000.fits, Empty file
wapp.20041116.a1961.0001.fits, bad data (trish_01)
wapp.20041116.a1961.0002.fits, Empty file
wapp.20041116.a1961.0003.fits
           (trish_01, trish_02) - adjust integration time to 2.5 min
wapp.20041116.a1961.0004.fits, Empty file
wapp.20041116.a1961.0005.fits, Empty file
wapp.20041116.a1961.0006.fits,
           (trish_01, trish_02) - adjust integration time to 2.7 min

wapp.20041116.a1961.0007.fits, Empty file
wapp.20041116.a1961.0008.fits,
           Large file. (trish_01 to trish_26)
           file was closed by cima, Too big!
           Interferences by puntas salinas for sources: trish_08 and trish_09
wapp.20041116.a1961.0009.fits, continuation of observations (trish_26 only)
wapp.20041116.a1961.0010.fits, observations of cal. 
(hcorcal)
Observations of ZOA source, from T. Henning.
Observers: Steve Schneider, Erik Muller, Robbie Auld

Startup failed - neglected to 'Adjust Power' this is an important step!
Use parallactic angle of 10.89

Tweaking of command scripts, too many ''newfile' commands

CIMA closed file before end of observing. We need to break up the file into useable data sizes.

The Cals did not appear to fire for as long as expected.

Using script observing, lfrog0_test.script,lfrog1_test.script,lfrog_all.v2.script
17th Nov
Using lfrog_all.v2.script
wapp.20041117.a1961.0000.fits, cals (hcorcal)
wapp.20041117.a1961.0002.fits to wapp.20041117.a1961.0029.fits
           (trish_01 to trish_28) Interference by L3
           on trish_12 and trish_13 at 1381MHz
wapp.20041117.a1961.0030.fits, empty file
wapp.20041117.a1961.0031.fits, observations of cal (hcorcal)
wapp.20041117.a1961.0032.fits, empty file
wapp.20041117.a1961.0033.fits to wapp.20041117.a1961.0036
           (trish_30 to trish_33), tweaked rotation parameters, still confused
Observations of ZOA source, from T. Henning.
Observers: Steve Schneider, Robbie Auld, Erik Muller

Tonight we want to test the alfa rotation, stability of multiple footprints.

Now observing with ONE pointing per datafile, this means that LIVEDATA does not operate!

Using script observing,lfrog_all.v2.script
18th Nov
Using lfrog_area_test.script (to fine-tune the integration times)
wapp.20041118.a1961.0001.fits, wapps got confused.
           Check 'apply backend setup' and repeat cal to see if this helps, seems ok.
wapp.20041118.a1961.0002.fits,
observations of cal (hcorcal)
wapp.20041118.a1961.0003.fits to wapp.20041118.a1961.0004.fits
           (lfrog1a_01, lfrog1a_04) - adjust integration time to 1.35 min
wapp.20041118.a1961.0005.fits to wapp.20041118.a1961.0006.fits
           (lfrog1a_01, lfrog1a_04) error in adjustment to script, fixed
wapp.20041118.a1961.0007.fits to wapp.20041118.a1961.0008.fits
           (lfrog1a_01, lfrog1a_04) - adjust integration time to 1.3 min
wapp.20041118.a1961.0009.fits to wapp.20041118.a1961.0010.fits
           (lfrog1a_01, lfrog1a_04) - adjust integration time to 1.25 min
           Should be good enough to continue. Confusion about rotation angle.
Using lfrog_area.script (the real script)
wapp.20041118.a1961.0011.fits, empty file
wapp.20041118.a1961.0012.fits to
wapp.20041118.a1961.0018.fits
           (lfrog1a_01,04,07,13,16,19,22)
           Error in catalog and script (missing pointing 10), timings are out of synch
           N.B., files 0012 0013 are contaminated by inteference from L3
wapp.20041118.a1961.0019.fits to wapp.20041118.a1961.0021.fits
           (lfrog1a_02,05,08)
           We are worried about the alfa rotation,
           change the parallactic angle to -10.89 and re-observe

wapp.20041118.a1961.0021.fits, empty file
wapp.20041118.a1961.0022.fits to wapp.20041118.a1961.0028.fits
           (lfrog1a_02,05,08,11,14,17,20)
           alfa angle =0, parallactic angle = -10.89, this setting seems to give expected results.
wapp.20041118.a1961.0029.fits to wapp.20041118.a1961.0035.fits
           (lfrog1a_03,06,09,13,15,18,21)
           alfa angle =0, parallactic angle = -10.89
wapp.20041118.a1961.0036.fits, empty file
wapp.20041118.a1961.0037.fits to wapp.20041118.a1961.0043.fits
           (lfrog1b_01,04,07,10,13,16,19)
           alfa angle =0, parallactic angle = -10.89
wapp.20041118.a1961.0044.fits, empty file
wapp.20041118.a1961.0045.fits to wapp.20041118.a1961.0051.fits
           (lfrog1b_02,05,08,11,14,17,20)
           alfa angle =0, parallactic angle = -10.89
wapp.20041118.a1961.0052.fits, empty file
Observations of scattered galaxy group at  +07 05 03.2  13 09 21.7 vel=8500.00 [Helio] to test the efficiency of interleaving pointings.
Observers: Steve Schneider, Erik Muller, Robbie Auld

We are still confused about the ALFA rotation parameter, positive or negative? When is the rotation updated?

It turns out that a rotate command is not sent to the array for each pointing control.  This means that we have to click and set the parallactic angle before each new pointing set. This is a bit of a hassle for now, but it should be fixable. We can  cope for the time being!


Using script observing, although because of problems with updating the array rotation, we have aborted the scripts to start a new set of leaps.
lfrog_area_test.script, lfrog_area.script


Post-observing comment:
We have decided that we must use a parallactic angle which is the negative of what we expected, this seems to give us the intended rotation of the array.
19th Nov
Using lfrog_area_p6.script
wapp.20041119.a1961.0000.fits, observations of cal (hcorcal)
wapp.20041119.a1961.0001.fits, empty file
wapp.20041119.a1961.0002.fits to wapp.20041119.a1961.0008.fits
           (lfrog1b_03,06,09,12,15,18,21)
Using lfrog_area_p1.script
wapp.20041119.a1961.0009.fits, observations of cal (hcorcal)
wapp.20041119.a1961.0010.fits to wapp.20041119.a1961.0016.fits
           (lfrog1a_01,04,07,10,13,16,19)
wapp.20041119.a1961.0017.fits, observations of cal (hcorcal)
Using zoa_B.script
wapp.20041119.a1961.0018.fits, observations of cal (hcorcal) (embedded in zoa_B_01)
wapp.20041119.a1961.0019.fits to wapp.20041119.a1961.0047.fits
          (zoa_B_01 until zoa_B_29)
           NOTE!: This data are contaminated by a wandering Lband birdy.
           Files 0031 to 0047 are affected (apparently the cause is known, but not the remedy).
wapp.20041119.a1961.0048.fits, observations of cal (hcorcal) (embedded in zoa_B_29)
         (N.B. This cal will be rubbish!, since the turret had already rotated)
Finishing Observations of galaxy group, then onto ZOA region, set B- having done set A on the 16th and 17th

Observers: Steve Schneider, Robbie Auld, Erik Muller

we have renamed the pointings here from trish_nn to zoa_nn

The process is becoming more automated, the outstanding problems are to
do with the rotation of the array. As yet, this is not updated after each
slew, nor is is something that can be put in the script.

Post observing note
, 20041124 - this data use 1400 as the rest freq
20th Nov
Using zoa_C.script
wapp.20041120.a1961.0000.fits, empty file
wapp.20041120.a1961.0001.fits, observations of cal (hcorcal) (embedded in zoa_C_01)
wapp.20041120.a1961.0002.fits to wapp.20041120.a1961.0038.fits
          (zoa_C_01 until zoa_C_37)
           L3 present in 0010, 0021 and 0022, Broadband rfi in 0017
wapp.20041120.a1961.0039.fits, empty file
wapp.20041120.a1961.0040.fits, observations of cal (hcorcal) (embedded in zoa_C_38)
Continuing ZOA region, set C, Robbie Driving.
Observers: Robbie Auld, Steve Schneider, Erik Muller

No problems observing.  interf from L3 appeared twice.
There was some asynchronis wapp start-times.
21st Nov
Using zoa_D.script
wapp.20041121.a1961.0000.fits, empty file
wapp.20041121.a1961.0001.fits, observations of cal (hcorcal) (embedded in zoa_D_01)
wapp.20041121.a1961.0002.fits to wapp.20041121.a1961.0033.fits
         (zoa_D_01 until zoa_D_32)
wapp.20041121.a1961.0035.fits, empty file
wapp.20041121.a1961.0035.fits, empty file
wapp.20041121.a1961.0002.fits to wapp.20041121.a1961.0039.fits
         (zoa_B_29 until zoa_B_32)
wapp.20041121.a1961.0040.fits, empty file
Continuing ZOA region, set D, SS driving.
No L3 tonight. woohoo (Trish deflected the Satellite, shot it down, sabotage, etc. etc. etc. )


Did some more of zoa_B
22nd Nov
Using zoa_fam1.script
wapp.20041122.a1961.0000.fits empty file
wapp.20041122.a1961.0001.fits observations of cal (hcorcal) (embedded in zoa_G_06)
wapp.20041121.a1961.0002.fits to wapp.20041121.a1961.0004.fits
Using zoa_fam2.script
wapp.20041122.a1961.0005.fits observations of cal (hcorcal) (embedded in zoa_I_06)
wapp.20041121.a1961.0006.fits to wapp.20041121.a1961.0008.fits
Using zoa_fam3.script
wapp.20041122.a1961.0009.fits observations of cal (hcorcal) (embedded in zoa_K_06)
wapp.20041121.a1961.0010.fits to wapp.20041121.a1961.0012.fits
Problems with calculated parallactic angle.
wapp.20041121.a1961.0013.fits to wapp.20041121.a1961.0015.fits
Recalculated angle is -22.02

MAY NEED TO REPEAT OBS BEFORE THIS TIME.

Using zoa_fam1.script

wapp.20041122.a1961.0016.fits empty file
wapp.20041122.a1961.0017.fits observations of cal (hcorcal) (embedded in zoa_G_06)
wapp.20041121.a1961.0018.fits to wapp.20041121.a1961.0020.fits
Using zoa_fam2.script
wapp.20041122.a1961.0021.fits observations of cal (hcorcal) (embedded in zoa_I_06)
wapp.20041121.a1961.0022.fits to wapp.20041121.a1961.0024.fits
Using zoa_fam3.script
wapp.20041122.a1961.0025.fits observations of cal (hcorcal) (embedded in zoa_K_06)
wapp.20041121.a1961.0026.fits to wapp.20041121.a1961.0028.fits
Using zoa_fam4.script
wapp.20041122.a1961.0029.fits observations of cal (hcorcal) (embedded in zoa_H_06)
wapp.20041121.a1961.0030.fits to wapp.20041121.a1961.0032.fits
Using zoa_fam5.script
wapp.20041122.a1961.0033.fits observations of cal (hcorcal) (embedded in zoa_J_06)
wapp.20041121.a1961.0034.fits to wapp.20041121.a1961.0036.fits
Using zoa_fam6.script
wapp.20041122.a1961.0037.fits observations of cal (hcorcal) (embedded in zoa_L_06)
wapp.20041121.a1961.0038.fits to wapp.20041121.a1961.0040.fits

Changed source to Trish's ZOA objects, using 'tmpscript.script' a modified
Using zoa_E.script start at a later position.
 
wapp.20041122.a1961.0041.fits observations of cal (hcorcal) (embedded in zoa_E_11)
Pointing seems odd for this cal
wapp.20041121.a1961.0042.fits
to
wapp.20041121.a1961.0064.fits
         (zoa_E_11 until zoa_E_34)
          L3 present in 0044, 0062, 0063
wapp.20041122.a1961.0065.fits empty file
Obeservers galore: Steve Schneider, Erik Muller, Robbie Auld, Trish Henning.

Post observing note, 20041124 - this data use 1400 as the rest freq
23rd Nov
Using script zoaN.script.
wapp.20041123.a1961.0000.fits to 0004  - empty file
wapp.20041124.a1961.0005.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041123.a1961.0006.fits to wapp.20041123.a1961.0010.fits
wapp.20041124.a1961.0011.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041123.a1961.0012.fits to wapp.20041123.a1961.0016.fits
wapp.20041124.a1961.0017.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041123.a1961.0018.fits to wapp.20041123.a1961.0022.fits
wapp.20041124.a1961.0023.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041123.a1961.0024.fits to wapp.20041123.a1961.0028.fits
wapp.20041124.a1961.0029.fits observations of cal (hcorcal)(embedded in first scan)
 
wapp.20041123.a1961.0030.fits to wapp.20041123.a1961.0034.fits
NB,  Cals written to 0030 as well as obs. data.
wapp.20041124.a1961.0035.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041123.a1961.0036.fits to wapp.20041123.a1961.0040.fits
Using script zoaS.script.
wapp.20041123.a1961.0041.fits empty file
wapp.20041124.a1961.0042.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041123.a1961.0043.fits to wapp.20041123.a1961.0048.fits
NB, error in script,, Cals written to 0043 as well as obs. data.
Using U405N.script
wapp.20041123.a1961.0048.fits
empty file
wapp.20041124.a1961.0049.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041123.a1961.0050.fits to wapp.20041123.a1961.0054.fits
wapp.20041124.a1961.0055.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041123.a1961.0056.fits to wapp.20041123.a1961.0060.fits
wapp.20041124.a1961.0061.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041123.a1961.0062.fits to wapp.20041123.a1961.0066.fits
wapp.20041123.a1961.0067.fits to wapp.20041123.a1961.0071.fits
NB, error in script,, Cals written to 0067 as well as obs. data.
wapp.20041124.a1961.0072.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041123.a1961.0073.fits to wapp.20041123.a1961.0077.fits
wapp.20041124.a1961.0078.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041123.a1961.0079.fits to wapp.20041123.a1961.0080.fits
          Ended part way through 0080
Observers: Trish H. Stephen S., Robbie A, Erik M.
Have arranged for 'PARAANGLE' to be a script keyword. Seems to be working well!

Some mis-calcs for the dec step of a Northerest strip to ZOA. A tepid drawback.

Post observing note, 20041124 - this data use 1400 as the rest freq
24th Nov
Using script zoaS.script.
wapp.20041124.a1961.0000.fits empty file

wapp.20041124.a1961.0001.fits to wapp.20041124.a1961.0002.fits
 Rubbish data, CIMA is not correctly configured (i.e. we forgot to source the new gui)
wapp.20041124.a1961.0003.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041124.a1961.0004.fits to wapp.20041124.a1961.0008.fits
wapp.20041124.a1961.0009.fits observations of cal (hcorcal)(embedded in first scan) wapp.20041124.a1961.00010.fits to wapp.20041124.a1961.0014.fits
wapp.20041124.a1961.0015.fits observations of cal (hcorcal)(embedded in first scan) wapp.20041124.a1961.0016.fits to wapp.20041124.a1961.0020.fits
wapp.20041124.a1961.0021.fits observations of cal (hcorcal)(embedded in first scan) wapp.20041124.a1961.0022.fits to wapp.20041124.a1961.0026.fits
    
Interference from L3 in 0025
wapp.20041124.a1961.0027.fits observations of cal (hcorcal)(embedded in first scan) wapp.20041124.a1961.0028.fits to wapp.20041124.a1961.0032.fits
wapp.20041124.a1961.0033.fits observations of cal (hcorcal)(embedded in first scan) wapp.20041124.a1961.0034.fits to wapp.20041124.a1961.0038.fits
wapp.20041124.a1961.0039.fits empty file
 
Using script U405S.script
wapp.20041124.a1961.0040.fits observations of cal (hcorcal)(embedded in first scan) wapp.20041124.a1961.0041.fits to wapp.20041124.a1961.0045.fits
wapp.20041124.a1961.0046.fits observations of cal (hcorcal)(embedded in first scan) wapp.20041124.a1961.0047.fits to wapp.20041124.a1961.0051.fits
wapp.20041124.a1961.0052.fits observations of cal (hcorcal)(embedded in first scan) wapp.20041124.a1961.0053.fits to wapp.20041124.a1961.0057.fits
wapp.20041124.a1961.0058.fits observations of cal (hcorcal)(embedded in first scan) wapp.20041124.a1961.0059.fits to wapp.20041124.a1961.0060.fits
wapp.20041124.a1961.0061.fits empty file
Observer: Trish, H. Robbie A, Stephen S, Erik M. AL. and..
The Frog of Destiny.

We have 11 beams!!.. it goes up to 11! - probably an error somewhere. Perhaps... no man, it goes all the way up to 11!


Post observing note, 20041124 - this data use 1400 as the rest freq
25th Nov
Using script zoaN2.script
wapp.20041125.a1961.0000.fits empty file
wapp.20041125.a1961.0001.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041124.a1961.0002.fits to wapp.20041124.a1961.0006.fits
wapp.20041125.a1961.0007.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041124.a1961.0008.fits to wapp.20041124.a1961.0012.fits
wapp.20041125.a1961.0013.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041124.a1961.0014.fits to wapp.20041124.a1961.0018.fits
wapp.20041125.a1961.0019.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041124.a1961.0020.fits to wapp.20041124.a1961.0024.fits
wapp.20041125.a1961.0025.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041124.a1961.0026.fits to wapp.20041124.a1961.0030.fits
wapp.20041125.a1961.0031.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041124.a1961.0032.fits to wapp.20041124.a1961.0036.fits
Using script u405X.script
wapp.20041125.a1961.0037.fits empty file
wapp.20041125.a1961.0038.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041124.a1961.0039.fits to wapp.20041124.a1961.0043.fits
wapp.20041125.a1961.0044.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041124.a1961.0045.fits to wapp.20041124.a1961.0049.fits
wapp.20041125.a1961.0050.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041124.a1961.0051.fits to wapp.20041124.a1961.0054.fits
wapp.20041125.a1961.0055.fits empty file
Observers: Robbie A. (Driving), Trish H.offering sage advice, Steve S. talking about beer, Bobbie M. listening to Steve, Erik M. making up biomass. AL on the laptop (again).

No interf. What a great dataset. Now if we could only reduce it....


There might be a problem with the script observing:
    1. We have to specify an IFLO setup, however this seems to cause errors
    2. When starting a new script, the 'starting line number' is not set to 0.

Due to .. um.. operator error, we missed the last set of leaps.

26th Nov
Using script zoaS2.script
wapp.20041126.a1961.0000.fits empty file

wapp.20041126.a1961.0001.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041126.a1961.0002.fits to wapp.20041126.a1961.0006.fits
wapp.20041126.a1961.0007.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041126.a1961.0008.fits to wapp.20041126.a1961.00012.fits
wapp.20041126.a1961.0013.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041126.a1961.0014.fits to wapp.20041126.a1961.0018.fits
wapp.20041126.a1961.0019.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041126.a1961.0020.fits to wapp.20041126.a1961.0024.fits
wapp.20041126.a1961.0025.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041126.a1961.0026.fits to wapp.20041126.a1961.0030.fits
wapp.20041126.a1961.0031.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041126.a1961.0032.fits to wapp.20041126.a1961.0036.fits
Using script zoa_E.script (note, this has been altered to include the PARAANG tag)
wapp.20041126.a1961.0037.fits empty file
wapp.20041126.a1961.0038.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041126.a1961.0039.fits to wapp.20041126.a1961.0048.fits
wapp.20041126.a1961.0049.fits empty file
completed E_01 to E_10
Observers: Robbie A. Driving again.. who lets that guy drive?, Erik M on the netscape and Steve Schneidy on the laptop.

Met the frog on the way in, had a ribbeting chat.
The mouse is buggered, no connection.


27th Nov
Using script zoa_F.script
wapp.20041127.a1961.0000.fits empty file
wapp.20041127.a1961.0001.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041127.a1961.0002.fits to wapp.20041127.a1961.0006.fits
wapp.20041127.a1961.0006.fits to
wapp.20041127.a1961.0009.fits - rubbish data.
wapp.20041127.a1961.0010.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041127.a1961.0011.fits to wapp.20041127.a1961.0032.fits
finished on zoa_F22
wapp.20041127.a1961.0033.fits - rubbish data.
Observers: Robbie A. being a telescope hog. Erik M, on the Netscape. Schneidy is whooping it up on the town.
No frogs tonight.

Something ghastly has happened to beams 1,3,5. The data are full of very small numbers. Of order 10-23!! (see what happens when there are no frogs?!)

Restarting wapps from pulsar control.
We got an 'unknown' file.

Ok, now appear to be getting spectra ok. Start again from F_01


28th Nov
Using script zoa_G.script
wapp.20041128.a1961.0000.fits empty file
wapp.20041128.a1961.0001.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041128.a1961.0002.fits to wapp.20041128.a1961.0033.fit
finished on G_32 now onto finish off F.
wapp.20041128.a1961.0034.fits to
wapp.20041128.a1961.0037.fits empty file
Using script tmp4script.script (for the last part of set F, start on F_23)
wapp.20041128.a1961.0038.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041128.a1961.0039.fits to wapp.20041128.a1961.0042.fit
Got turfed from telescope with 40 sec remaining to go on last scan! where is the love?!
- last complete scan on F_27 - just  enough.
N.B. file
wapp.20041128.a1961.0043.fit does not contain a full scan of F_28.
Observers: Sabina S. Driving, Robbie A. lecturing, Erik M. on the netscape, Bobbie as biomass. Sans Steve S.

No frogs. Not even an ugl
y toad.

AZ is about 360 deg for G
29th Nov
Using script zoa_H.script
wapp.20041129.a1961.0000.fits empty file
wapp.20041129.a1961.0001.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041129.a1961.0002.fits (dud data..)
Aborted, didn't source the correct gui code..
wapp.20041129.a1961.0003.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041129.a1961.0004.fits  to wapp.20041129.a1961.0035.fit
last complete scan on H_32
wapp.20041130.a1961.0036.fits empty file
Using script zoa_F_tail.script
wapp.20041129.a1961.0037.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041129.a1961.0038.fits to wapp.20041129.a1961.0042.fit
last complete scan on F_32
Observers,  Sabina S. Driving, Robbie A lecturing, Erik M on the netscape, Steve S. as eeevil overlord.

AZ is about 348 deg  for H.

Are we seeing some kind of position error in the beams?
wapp.20041129.a1961.0004.fits looks as there is some jitter.
30th Nov
Using script zoa_I.script
wapp.20041130.a1961.0000.fits empty file
wapp.20041130.a1961.0001.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041130.a1961.0002.fits  to wapp.20041130.a1961.0033.fits
last complete scan on I_32
wapp.20041130.a1961.0034.fits empty file
Using script zoa_end.script
wapp.20041130.a1961.0035.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041130.a1961.0036.fits  to wapp.20041130.a1961.0041.fits
wapp.20041130.a1961.0042.fits empty file
Observers: Dr. Sabtini Driving, Robbie A. Irritator-in-cheif, Erik M. on netscape. Steve S. eeevil overlord extrodinaire.
NB. The Cryos failed earlier today, probably this data is tripe.
Looking at the power from the cal, and at the rms's for today and yesterday gives similar levels.... who knows? not I. We will bravely continue bravely.

AZ is about 353 deg  for I

AZ is about 387 for zoa_end

1st Dec

Using script zoa_J.script
wapp.20041201.a1961.0000.fits empty file
wapp.20041201.a1961.0001.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041201.a1961.0002.fits  to wapp.20041201.a1961.0033.fits
Using script zoa_end.script
wapp.20041201.a1961.0035.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041201.a1961.0036.fits  to wapp.20041201.a1961.0041.fits
wapp.20041201.a1961.0042.fits empty file

White Frogs!

Observers: Dr. Steve Schneider captaining, Robbie A. (proto Dr.)
no. 1. Absentees: Dr. Erik Muller, Dr. Sabina Sabatini (in bed... not the same bed obviously) Dr. Robert 'alfa_ang' Minchin making a late appearance, procrastinating when he should be writing his talk no doubt. Willy depressing us with the state of the Peurto Rico streets, neighbourhoods, drug problems etc.

RFI: Wandering birdie  in 0004, 0005, 0011, 0012, 0013, 0017, 0018, 0019, 0023, 0024, 0028, 0029, 0030, 0033, 0034

AZ is about 359 deg for J
L3 in 0020

Messrs. Schneider and Auld would like to report a bug in pos_mon whereby a colour change is happening every five pointings. Handy for highlighting families but NOT for long runs at the same dec. Pretty though.

This is Robbie A. signing off for this observing run.

2nd Dec
Using script zoa_K.script
wapp.20041202.a1961.0000.fits empty file
wapp.20041202.a1961.0001.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041202.a1961.0002.fits  to wapp.20041202.a1961.0033.fits
last complete scan on K_32
wapp.20041202.a1961.0034.fits empty file

Using script zoa_end.script
wapp.20041202.a1961.0035.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041202.a1961.0036.fits  to wapp.20041201.a1961.0042.fits
We have some time left over!, do zoa_end again!
wapp.20041202.a1961.0043.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041202.a1961.0044.fits  to wapp.20041202.a1961.00.fits
wapp.20041202.a1961.0046.fits empty file
Observers, Dr. Sabatini Driving, Eeevil Overlord Schneider presiding. Erik M. on the keyboard.

AZ is about 348 deg for K.
AZ is about 383 for end (for first loop)
AZ is about 408 for end (for second loop)

RFI:
Wandering birdie  in 0008 (maybe earlier) and onwards. Forever. All data is stuffed.
Also: Broadband rfi in 0031
Also: short, 1min hit of L3 in 0044

Mr. Muller would like to report the satisfactory completion of the new efforts which have been made to oblige the previous nights request regading the pos_mon facility, while preseving the inherent prettiness. By the way: it's not a bug, it was deliberate. ner.

3rd Dec
Using script zoa_L.script
wapp.20041203.a1961.0000.fits empty file
wapp.20041203.a1961.0001.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041203.a1961.0002.fits  to wapp.20041203.a1961.0033.fits
last complete scan on L_32
wapp.20041203.a1961.0034.fits empty file

Using script zoa_end.script
wapp.20041203.a1961.0035.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041203.a1961.0036.fits  to wapp.20041203.a1961.0041.fits
wapp.20041203.a1961.0042.fits empty file
Observers: Eevil overlord SS driving. Semi-evil minon EM on keyboard, SS also as moral support. 'Weevil' Overlord supreme JD in training (WOJD and SSII couldn't take the heat, piked out early).

The Wandering birdy is wandering all over our data.... again.
Also, the L3 threw up in our data. (EOSS's words). Affecting 0027
 
4th Dec
Using script mess.script
wapp.20041204.a1961.0000.fits empty file

wapp.20041204.a1961.0001.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041204.a1961.0002.fits to wapp.20041204.a1961.0033.fits wapp.20041204.a1961.0034.fits empty file
Observers: EOSS is driving, Minion EM on netscape.
Beam 5 appears to be schnookered. It isnt working.
We have tried to restart the WAPPS, to no avail.
We have the puntas selinas radar at 1367 and 1382 MHz
The Frog god smiles on us no longer.
5th Dec
Using script muck.script
wapp.20041205.a1961.0000.fits empty file
wapp.20041205.a1961.0001.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041205.a1961.0002.fits to wapp.20041204.a1961.0030.fits
wapp.20041205.a1961.0031.fits empty file
Observers: SS (jr) Driving the metal behemoth. EM sans laundry.
SS (sr) and WVD out partying up on the town, but are here in spirit nonetheless.
 
Birdy ALL OVER THE DATA!!!! -
yes, again.. will somone kill it?! PLEASE?
specifically: files 0001-0029
6th Dec Using script zoa_A.script
wapp.20041206.a1961.0000.fits empty file
wapp.20041206.a1961.0001.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041206.a1961.0002.fits to wapp.20041204.a1961.0007.fits
wapp.20041205.a1961.0008.fits empty file
Using script zoa_end.script
wapp.20041206.a1961.0009.fits observations of cal (hcorcal)(embedded in first scan)
wapp.20041206.a1961.0010.fits to wapp.20041204.a1961.0015.fits
wapp.20041205.a1961.0016.fits empty file
Observers: SS (sr - EOL) , driving, and WvD (as himself).

Soaring freely through
water falling in sea-blue:
Birdy all askew.
(Haiku by Song Sung Lo, 1340-1420).

ZOA table o' progress 
(Green indicates data taken using incorrect parallactic angle,
purple indicates data taken with the correct parallactic angle)


1
2
3
4
5
6
7
8
9
10
11
12
13
14
15
16
17
18
19
20
21
22
23
24
25
26
27
28
29
30
31
32
A
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
B
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
C
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
D
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
E
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
F
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
G
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
H
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
I
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
J
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1 1
1
1
1
1
1
1
1
1
1
1
1
1
1
K
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
L
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
1
end

























6
6
5
5
5
5
1